# Open-vocabulary Multimodal Emotion Recognition: Dataset, Metric, and Benchmark

### v1

@@ Appendix A Related Work
Appendix B More Examples
In Figures 9$\sim$11, we provide more examples to visualize the difference between one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. We use DemoAI to remove personal information.
@@ Appendix B More Examples
Appendix C Summary of Abbreviations
@@
Appendix D Details in Dataset Construction
Figure 2 presents our dataset construction process. In Table 6, we provide prompts and corresponding models used in this process. In Figure 12, we visualize the output of the main steps.
Function (Model) Prompt
[…]
Figure 12: An example to visualize the output of the main steps.
Appendix E One-hot vs. OV Labels
(a) Word cloud
(b) Label number distribution
Figure 13: Word cloud and label number distribution of OV labels.
Table 7: Performance of one-hot labels in OV-MER.
Language $\mbox{F}_{\mbox{s}}\uparrow$ $\mbox{Precision}_{\mbox{s}}\uparrow$ $\mbox{Recall}_{\mbox{s}}\uparrow$
English 65.71±0.06 92.17±0.00 51.05±0.08
Chinese 66.16±0.02 93.07±0.00 51.32±0.03
Figure 14: Emotion distribution of OV labels.
Appendix F Dataset Comparison
This paper introduces a new task, OV-MER, and constructs a dataset for this task called OV-MERD. Table 8 compares OV-MERD with existing datasets. The annotation types of these datasets can be broadly categorized into two types: dimensional emotions and discrete emotions. We classify sentiment analysis datasets (e.g., CMU-MOSI) as dimensional datasets because the definition of sentiment intensity overlaps with the valence in dimensional emotions. We observe that OV-MERD contains 248 emotion categories, with most samples having 2 to 4 labels, significantly exceeding the number of labels in existing datasets. In the future, as the scale of the dataset increases, the number of candidate labels can be further expanded.
Table 8: Dataset comparison. “I”, “A”, “V”, and “T” are abbreviations for image, audio, video, and text, respectively.
@@ Appendix F Dataset Comparison
OV-MERD (Ours) A,V,T Discrete 248
@@
Appendix G Annotation Details
@@
Figure 15 shows the layout of the annotation platform used for manually checking acoustic and visual clues. During annotation, we use the following instructions: We provide pre-labeled acoustic and visual clues. Please manually check these clues, remove errors, and add missing information. To reduce the difficulty of annotation, we design an interface with a time slider, allowing annotators to start playing the video from any frame to capture subtle changes within the video.
@@
Appendix H Jaccard Similarity Coefficient
The motivation for the above metrics is that $\mathcal{Y}$ represents the ground truth, while $\hat{\mathcal{Y}}$ represents the prediction. However, in Figure 2, the two sets of emotions are considered equally important. Therefore, we use the Jaccard similarity coefficient to measure the similarity. This metric evaluates the similarity between two sets by comparing the size of their intersection to the size of their union:
Appendix I CLUE-Multi Analysis
@@
Appendix J Visualization of CLUE-M/A/T/V
@@
Appendix K Details of CLUE-MLLM
@@
Appendix L GPT-based vs. Matching-based Metrics
@@
Appendix M Emotion Wheel
[…]
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5

### v2

@@ Appendix D Related Work
Appendix E More Examples
Figures 8$\sim$10 provide more examples to visualize the difference between one-hot and OV labels. This paper uses emotion-related descriptions as a bridge for OV label extraction. Since the original video contains real people, we use DemoAI to remove personal information to address copyright concerns. Our OV-MERD dataset is derived from the MER2023 dataset [9] with further annotations. Therefore, for the original data, please download the MER2023 dataset.
@@ Appendix E More Examples
Appendix F Summary of Abbreviations
@@
Appendix G Dataset Comparison
This paper introduces a new task, OV-MER, and constructs a dataset for this task called OV-MERD. Table 5 compares OV-MERD with existing datasets. The annotation types of these datasets can be broadly categorized into two types: dimensional emotions and discrete emotions. We classify sentiment analysis datasets (e.g., CMU-MOSI) as dimensional datasets because the definition of sentiment intensity overlaps with the valence in dimensional emotions. We observe that OV-MERD contains 236 emotion categories, with most samples having 2 to 4 labels, significantly exceeding the number of labels in existing datasets. In the future, as the scale of the dataset increases, the number of candidate labels can be further expanded. Meanwhile, we would like to emphasize that our OV-MERD is the first dataset that uses the human-LLM collaborative annotation strategy, aiming to provide richer labels to capture more nuanced emotions. We believe this work is an important extension of traditional MER and will contribute to the development of the field.
Table 5: Dataset comparison. In this table, “I”, “A”, “V”, and “T” are abbreviations for image, audio, video, and text, respectively. Some datasets (such as CMU-MOSEI and MSP-Podcast) contain both discrete and dimensional emotions.
@@ Appendix G Dataset Comparison
MSP-Podcast [42] A Discrete 8 1
JL-Corpus [52] A Discrete 10 1
EmoDB [53] A Discrete 7 1
EMOVO [54] A Discrete 7 1
MESD [55] A Discrete 6 1
@@
RAF-DB [7] I Discrete 19 1$\sim$2
CMU-MOSEI [48] A,V,T Discrete 6 1
eNTERFACE [61] A,V,T Discrete 6 1
SAVEE [62] A,V,T Discrete 7 1
@@
OV-MERD (Ours) A,V,T Discrete 236
@@
Appendix H One-hot vs. OV Labels
(a) Word cloud
(b) Label number distribution
Figure 11: Word cloud and label number distribution of OV labels.
Table 6: Performance of one-hot labels in OV-MER.
Language $\mbox{F}_{\mbox{s}}$uparrow $\mbox{Precision}_{\mbox{s}}$uparrow $\mbox{Recall}_{\mbox{s}}$uparrow
English 65.71±0.06 92.17±0.00 51.05±0.08
Chinese 66.16±0.02 93.07±0.00 51.32±0.03
Figure 12: Emotion distribution of OV labels.
Appendix I Duration Distribution of OV-MERD
In Figure 13, we analyze the duration distribution of the OV-MERD dataset. We observe that the majority of the samples have durations ranging from 1 to 4 seconds. This distribution is consistent with that of the MER2023 dataset, which was used as the original dataset for constructing OV-MERD (see Section 2.3 for details).
Figure 13: Duration distribution of the OV-MERD dataset.
Appendix J Details in Dataset Construction
Prompts.
Figure 2 presents our dataset construction process. In Table 7, we provide prompts and corresponding models used in this process.
Function (Model) Prompt
[…]
Number of Sampled Frames in Pre-annotation
To generate pre-annotated visual clues, we sample three frames from each video and input them into GPT-4V. In this section, we discuss the rationale behind the choice of the number of sampled frames. Specifically, we categorize visual clues into two types: (1) visual clues with relatively long durations; and (2) visual clues with fast movements, such as eye movements, head movements, and micro-expressions. For the first type, since the duration of most videos is between 1 and 4 seconds (see Appendix I) and the video content is usually continuous with only minor differences between adjacent frames (see Appendix E), uniformly sampling three frames are sufficient to capture this information; For the second type, we observe that current MLLMs (including the GPT-4V used in this paper) struggle to capture these fast movements. Increasing the number of sampled frames does not address this issue. Previous research has also shown that GPT-4V cannot recognize micro-expressions [69]. To capture these fast movements, we employ multiple professional annotators to manually add this information.
Merging Process.
In this paper, we rely on the powerful reasoning capabilities of LLM for multimodal fusion. Specifically, as shown in Table 7, we ask LLM to integrate textual, acoustic, and visual clues to infer the emotional state. From the output, we observe that LLM can produce reasonable analytical results. However, ambiguities and contradictions in multimodal fusion are inevitable, which is a challenging and open problem. In this paper, we simply use LLM to address this problem, providing a practical solution for the OV-MER task. However, more effective strategies may exist, such as using more complex prompts or incorporating modality control measures. Therefore, we plan to explore this area in the future.
Visualization.
In Figure 14, we visualize the output of the main steps.
Figure 14: An example to visualize the output of the main steps.
Appendix K Annotation Details
@@ Appendix I Duration Distribution of OV-MERD
Figure 15 shows the layout of the annotation platform used for manually checking acoustic and visual clues. During the annotation process, we use the following instructions: We provide pre-labeled acoustic and visual clues. Please manually check these clues, remove errors, and add missing information. On the annotation platform, we design an interface with a time slider, allowing annotators to start playing the video from any frame. This enables annotators to view the entire video during the manual check, helping them better annotate the details that may be missed in pre-annotation.
@@
To illustrate which labels are removed or kept, we provide two examples, each with labels from multiple annotators. To ensure annotation quality, we hire professional annotators who are experts in affective computing and familiar with the definition of emotions. Some of these annotators are members of our team who specialize in affective computing. In Figure 17, as the character doesn’t know which doctor to see, most annotators provide labels such as confused or puzzled. Based on his tone and expression, some annotators further provide labels like anxious and serious. In Figure 18, most annotators notice his disapprove based on the textual content. Combining other modalities, some annotators further note his blame and accuse of what others are planning to do. From these examples, we can observe that these annotators provided relatively reliable labels. However, some annotators may focus only on the most relevant labels and overlook some details. To ensure the comprehensiveness of the annotation results, we merge the labels checked by four annotators. For example, in Figure 17, the final merged labels are troubled, focused, puzzled, anxious, worried, confused, and serious. In the next round, we invite another four annotators for a second check. Through this process, we can ensure that each preserved label is confirmed by at least one annotator in each round, thereby ensuring the comprehensiveness and accuracy of annotation results.
Figure 17: Example1 with labels from multiple annotators.
Figure 18: Example2 with labels from multiple annotators.
Appendix L CLUE-Multi Analysis
@@
Appendix M Details of Language Impact Experiments
Experimental Design.
In Figure 2, we analyze from two perspectives: 1) the impact of descriptive language (Clue-Multi), and 2) the impact of abstract language (OV labels). The $Y_{\text{EE}}$ to $Y_{\text{EC}}$ (or $Y_{\text{CC}}$ to $Y_{\text{CE}}$) experiment aims to keep the descriptive language consistent to analyze the effect of abstract language, while the $Y_{\text{CE}}$ to $Y_{\text{EE}}$ (or $Y_{\text{EC}}$ to $Y_{\text{CC}}$) experiment aims to keep the abstract language consistent to analyze the effect of descriptive language.
Jaccard Similarity Coefficient.
The motivation for the above metrics is that $\mathcal{Y}$ represents the ground truth, while $\hat{\mathcal{Y}}$ represents the prediction. However, in Figure 2, the two sets of emotions are considered equally important. As a result, we use the Jaccard similarity coefficient to measure the similarity. This metric evaluates the similarity between two sets by comparing the size of their intersection to the size of their union:
Appendix N Cost of GPT-based Metrics
This paper reports zero-shot performance, focusing only on the inference process. The cost of evaluating our OV-MERD dataset is approximately \$1 per evaluation, which may not seem high. However, for future work aimed at training frameworks to better address the OV-MER task, this cost will become prohibitive. For example, if we plan to train a model for 100 epochs, the evaluation cost will rise to \$$\$1\times 100\;\text{epochs}=\$100$. If we intend to test $N$ different parameter combinations and $M$ different frameworks, the evaluation cost will increase to \$$\$100\times M\times N$. Moreover, we plan to expand the OV-MERD dataset in the future. This cost will further increase. Therefore, this paper explores alternatives to GPT-based metrics.
Appendix O Emotion Wheel
[…]
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5
Appendix P Visualization of CLUE-M/A/T/V
@@
Appendix Q Details of CLUE-MLLM
@@
Appendix R Relationship between Description Length and Label Numbers
This section further discusses the relationship between description length and the number of labels per sample, i.e., whether longer descriptions correlate with more labels. To this end, we compute their PCC scores. We observe that, for the human-only strategy, the PCC score is 0.3416, and for the human-LLM collaboration strategy, the PCC score is 0.2939. Therefore, although from the dataset level, the length of descriptions is related to the richness of labels (see Figure 5), these two metrics do not show a strong correlation at the sample level.
Appendix S GPT-based vs. EW-based Grouping
We explore the relationship between GPT-based and EW-based grouping strategies. We calculate the scores for different types of EW-based grouping strategies, with the average score denoted as M-avg. Table 12 reports $\mbox{F}_{\mbox{s}}$, as this metric is used for the final ranking, and we also compute the PCC scores between different metrics. We observe that the PCC score between M-avg and GPT is relatively high, indicating that EW-based metrics can serve as alternatives to GPT-based metrics.
Model GPT M1 M2 M3-W1 M3-W2
L1 L2 L1 L2
PCC score 0.887 0.857 0.911 0.940 0.913 0.942
Model M3-W3 M3-W4 M3-W5 M-avg
L1 L2 L1 L2 L1 L2
PCC score 0.904 0.927 0.899 0.922 0.885 0.894 0.942
Appendix T GPT-based vs. Matching-based Metrics
@@
In Table 13, we observe that there is no strong correlation between the GPT-based metrics and the matching-based metrics. To clarify this point, we use the following three sentences as examples:
#1. The clue is “the weather is great”. His emotion is “happy”.
#2. The clue is “the weather is bad”. His emotion is “sad”.
#3. His emotion is “happy”.
For matching-based metrics, we use BLEU1 as an example. The BLEU1 score between #1 and #2 is 0.8181, while the BLEU1 score between #1 and #3 is 0.1738. Therefore, based on the BLEU1 score, #1 is closer to #2. For LLM-based metrics, we first extract the emotion labels and compare their similarity, so #1 is closer to #3. This demonstrates that matching-based metrics are not suitable for evaluating emotion recognition performance.